OpenStudy (anonymous):

Solve by elimination, equivalent system or substitution. x+6z=12 -2x+3y=6 y-z/2=5/2

OpenStudy (anonymous):

rearrange the first one in terms of x or z first.

OpenStudy (anonymous):

x=-6z+12 -2x+3y=6 -2(-6z+12)+3y=6 12z-24+3y=6 3y+12z=30 y=-4z+10 -4z+10-z/2=5/2 -4z-z/2=-7.5 Do I divide the 2 to both sides?

OpenStudy (anonymous):

\[\frac{-4z+10-z}{2} =\frac{5}{2}\] \[-5z + 10 = \frac{5}{2}\times 2\] -5z + 10 = 5 -5z = 5-10 -5z=-5 z=-5/-5 z=1

OpenStudy (anonymous):

x+6z=12 x+6(1)=12 x=6 -2+3y=6 -2(6)+3y=6 -12+3y=6 3y=18 y=6 Solution: (6,6,1)

OpenStudy (anonymous):

Am I correct?